# thanks: Jointly supervised this workthanks: Jointly supervised this work

### v1

@@ Abstract
Dissipation fundamentally alters quantum transport in Josephson junctions. Here, we demonstrate unique transport signatures in dissipative topological insulator-based Josephson junctions, which provide a powerful platform for exploring the competition between dissipation and topological protection. We incorporate dissipation effects by coupling a ‘lossy’ metallic lead at the junction to an electron reservoir, and describe its effects using an effective non-Hermitian Hamiltonian within the Lindblad formalism. Non-Hermiticity introduces finite lifetimes to the helical Andreev bound states in the junction, with imaginary energy components that depend on the quasiparticle incidence angle, while Klein tunneling remains robust for normal incidence. Unlike ‘ordinary’ non-Hermitian Josephson junctions, this system exhibits no Josephson gaps or forbidden regions, preserving topological protection. In the fully non-Hermitian regime, a tunable line of zero-energy states emerges, bounded by exceptional-like points with divergent supercurrents. Our results reveal how dissipation reshapes superconducting transport in topological insulator-based Josephson junctions, opening new directions for dissipation-engineered hybrid quantum devices.
@@
Josephson junctions (JJs) play a pivotal role in superconducting electronics, offering a versatile tool for probing quantum transport, topological states, and unique superconducting behaviors [1]. In conventional Hermitian JJs, the superconducting phase difference determines the current-phase relationship (CPR), which is typically $2\pi$-periodic and sinusoidal in ‘short’ junctions [2]. These junctions are fundamental to technologies such as Superconducting Quantum Interference Devices (SQUIDs) and single-photon detectors [3]. Recent advances have revealed non-reciprocal effects, such as the Josephson diode effect, in junctions with broken inversion and time-reversal symmetries [4, 5, 6, 7]. Topological insulator-based Josephson junctions (TIJJs) provide an ideal setting for realizing exotic quantum states [8, 9, 10, 11], where proximity-induced superconductivity on the surface of a 3D topological insulator (3DTI) induces helical Andreev bound states (ABSs) [8, 12, 13]. These junctions support non-chiral Majorana modes [8, 14, 15], leading to $4\pi$-periodic Josephson effects [8, 16, 13, 17].
Dissipation is an intrinsic aspect of real-world Josephson junctions, where non-Hermitian Hamiltonians capture the effects of environmental coupling [18, 19, 20, 21, 22, 23]. In non-Hermitian Josephson junctions (NH-JJs), dissipation introduces complex eigenspectra and exceptional points, leading to fundamentally different behavior from their Hermitian counterparts [20, 18, 21]. NH-JJs coupled to fermionic environments can develop phase intervals devoid of ABSs, known as Josephson gaps [21, 18]. Additionally, non-Hermitian effects modify the Josephson current and lead to novel superconducting transport phenomena [20, 18]. However, the impact of non-Hermiticity on topological Josephson junctions remains largely unexplored, particularly in systems where dissipation competes with topological protection.
@@
Figure 2: Characteristics of the NH-TIJJ for a general (complex) barrier ($Z_{1}\neq 0,Z_{2}\neq 0$): (a) The real (solid-lines) and imaginary (dashed-lines) parts of the ABS energy spectrum as a function of the phase difference $\phi$ for different values of the angle of incidence $\theta$ taking the imaginary potential barrier strength $Z_{2}=0.1$. (b) The variation of the energy gap in the real part of the ABS energy spectrum ($E_{\rm gap}=\text{Re}[E^{(+)}_{b}-E^{(-)}_{b}]$) at $\phi=\pi$. (c) The real (solid-lines) and imaginary (dashed-lines) parts of the angle-resolved supercurrent $I(\theta,\phi)$ plotted as a function of $\phi$ for different $\theta$. (d) The total supercurrent $I_{\rm tot}(\phi)$ as a function of $\phi$. We have chosen the real potential barrier strength $Z_{1}=0.4$ and the tunneling parameter $\gamma=0.3$.
In this work, we address the central question: how does dissipation caused by coupling a ‘lossy’ metallic lead [18, 24] at the junction to an electron reservoir, affect the quantum transport properties of TIJJs? To explore this, we model the dissipative effects of the lead by an effective non-Hermitian Hamiltonian [25, 20] and investigate transport in a non-Hermitian lateral Josephson junction on the surface of a 3DTI (NH-TIJJs), as illustrated schematically in Fig. 1. Here, proximity-induced superconductivity intertwines with non-Hermitian effects, offering a unique platform to study the interplay between dissipation and topological superconductivity.
Effective model of a NH-TIJJ: The NH-TIJJ in Fig. 1 consists of two $s$-wave singlet superconductors with identical gap $\Delta$ and phases $\phi_{L}$ and $\phi_{R}$ on the left and right, respectively. Dissipation due to the ‘lossy’ lead [18, 24] connected to an electron reservoir is modeled in a minimal way, within the Lindblad formalism [26, 27] (see Appendix A for details) and neglecting quantum jumps, by an imaginary potential barrier at the junction [20]. The combined effect of the top gate and the lead is therefore captured by a complex contact junction potential barrier: $U(x)=(V_{1}-iV_{2})\delta(x)$, where $V_{1}>0$ represents the real barrier strength tunable by the applied gate voltage $V_{g}$ and $V_{2}>0$ accounts for the imaginary barrier strength induced by the lead.
Assuming tunnel coupling between the superconductors and the 3DTI surface state [12, 8, 28, 29], the physics of the proximity-induced superconductivity on the surface state can be described by the following effective Bogoliubov-de Gennes (BdG) Hamiltonian [12, 8]:
@@
$H_{\rm BdG}$ $=$ $\begin{bmatrix}\frac{\hbar}{i}v_{F}\bm{\sigma}\cdot\nabla-\mu-\Gamma(\epsilon)&\tilde{\Delta}(\epsilon)i\sigma_{y}e^{i\phi_{L,R}}\vspace{0.1cm}\\ -\tilde{\Delta}(\epsilon)i\sigma_{y}e^{-i\phi_{L,R}}&\frac{\hbar}{i}v_{F}\bm{\sigma}^{*}\cdot\nabla+\mu-\Gamma(\epsilon)\end{bmatrix},$
$\tilde{\Delta}(\epsilon)$ $=$ $\frac{i\Gamma_{0}\Delta}{\sqrt{\epsilon^{2}-\Delta^{2}}}\;,\Gamma(\epsilon)=\frac{i\Gamma_{0}\epsilon}{\sqrt{\epsilon^{2}-\Delta^{2}}}\;,\Gamma_{0}=\pi t^{2}\mathcal{D}_{s}.$ (1)
@@
Here, $\Phi^{\dagger}=(\varphi^{\dagger}_{\uparrow},\varphi^{\dagger}_{\downarrow},\varphi_{\uparrow},\varphi_{\downarrow})$ is the Nambu basis and $\varphi^{\dagger}_{s}$ is the electron creation operator with spin $s=\uparrow,\downarrow$. $\mu$ is the chemical potential, $v_{F}$ is the Fermi velocity of the surface states of the 3DTI and $\sigma_{i}$ ($i=1,2$) are the Pauli matrices in spin space. $\tilde{\Delta}(\epsilon)$ is the proximity induced superconducting pairing potential which is a function of the Bogoliubov quasi-particle energy $\epsilon$, $\Gamma(\epsilon)$ is the shift in energy of the quasi-particles due to tunneling with amplitude $t$, and $\Gamma_{0}$ is the tunneling energy scale that determines the tunneling rate. $\mathcal{D}_{s}$ is the density of states of the superconductors in the normal state.
Figure 3: Characteristics of the NH-TIJJ for a purely dissipative barrier ($Z_{1}=0,Z_{2}\neq 0$): (a) The real (solid-lines) and imaginary (dashed-lines) parts of $E_{b}$ as a function of $\phi$ for different $\theta$ taking $Z_{2}=0.1$. (b) The real part of the ‘+’ branch of the ABS energy spectrum at $\theta=\pi/3$. The blue region shows zero-energy states (ZES) and the change in its boundary demonstrates tuning of the line of ZES (dotted line corresponding to the analytical expression in Eq. (9)) by the non-Hermiticity strength $Z_{2}$. (c) The real (solid-lines) and imaginary (dashed-lines) parts of $I(\theta,\phi)$ as a function of $\phi$ for different $\theta$. (d) Total supercurrent $I_{\rm tot}(\phi)$ as a function of $\phi$ for the fully transparent ($Z_{2}$=0) and purely non-Hermitian ($Z_{2}$=0.1) cases.
@@
To investigate transport in the ‘short’ NH-TIJJ (Fig. 1), assuming the width of the junction much smaller than the coherence length of the superconductors, we consider $\phi_{L}=0$ and $\phi_{R}=\phi$ without loss of generality. In this limit, the physics of the system can be described by [12]:
$\left[H_{\text{BdG}}(\bm{r})+\tau_{z}\sigma_{0}U(x)\right]\Psi(\bm{r})=\epsilon\Psi(\bm{r})\;,$ (2)
@@
where, $\tau_{z}$ is the third Pauli matrix in the Nambu space, $\sigma_{0}$ is the $2\times 2$ identity matrix and $\Psi(\bm{r})=\left[\psi_{\uparrow,\epsilon}(\bm{r}),\psi_{\downarrow,\epsilon}(\bm{r}),\psi^{*}_{\uparrow,-\epsilon}(\bm{r}),\psi^{*}_{\downarrow,-\epsilon}(\bm{r})\right]^{T}$ is the Nambu spinor wavefunction. Quasiparticles enter the barrier from the left at an angle of incidence $\theta\in[-\pi/2,\pi/2]$ and we assume translational invariance along the $y$-direction. The exact shape of the barrier potential is unimportant here, as our focus is on the effects of non-Hermiticity on Klein tunneling, which remains unaffected with the change in the barrier’s form [12, 30].
@@
$\psi_{k_{y}}(0_{-})=[1+iZ\tau_{z}\sigma_{x}]\psi_{k_{y}}(0_{+})\;,$ (3)
which leads to an equation for $\alpha=\left|\frac{\epsilon}{\Delta}\right|^{2}$:
@@
$\alpha\left(1-\alpha\right)=\gamma^{2}\left(\sqrt{1-\mathcal{\zeta}(Z,\theta)\sin^{2}\left(\frac{\phi}{2}\right)}-\alpha\right)^{2}\,.$ (4)

### v2

@@ Abstract
The interplay between topological protection and dissipation constitutes a critical frontier in the realization of hybrid quantum devices. Here, we investigate the transport signatures in a dissipative topological insulator-based Josephson junction, a platform that directly probes the competition between quantum coherence and loss. We model dissipation by coupling a ‘lossy’ metallic lead to the junction, described effectively by a non-Hermitian Hamiltonian derived using the Lindblad formalism. We observe that the junction exhibits an asymmetric complex Andreev spectrum, where the imaginary energy component imposes a finite lifetime on the quasi-bound states. Furthermore, beyond specific phase intervals, the real component of the spectrum bifurcates: one branch merges with the continuum, while the other penetrates just below the superconducting gap. Crucially, the characteristic zero-energy crossing shifts away from $\phi=\pi$ and acquires a non-zero imaginary component; consequently, the associated Majorana bound states acquire a finite lifetime, signaling a loss of robustness against dissipation. Finally, this spectral asymmetry drives an anomalous supercurrent, manifested as a non-vanishing current at zero phase difference. Our results reveal how dissipation fundamentally reshapes superconducting transport in topological junctions, opening new directions for dissipation-engineered quantum devices.
@@
Josephson junctions (JJs) play a pivotal role in superconducting electronics, offering a versatile tool for probing quantum transport, topological states, and unique superconducting behaviors [Wolf2017]. In conventional Hermitian JJs, the superconducting phase difference determines the current-phase relationship (CPR), which is typically $2\pi$-periodic and sinusoidal in ‘short’ junctions [Josephson1974]. These junctions are fundamental to technologies such as Superconducting Quantum Interference Devices (SQUIDs) and single-photon detectors [walsh2021]. Topological insulator-based Josephson junctions (TIJJs) provide an ideal setting for realizing exotic quantum states [Fu2008, Cayao2022a, Dutta2023, Dutta2024], where proximity-induced superconductivity on the surface of a 3D topological insulator (3DTI) induces helical Andreev bound states (ABSs) [Fu2008, Tkachov2013, Olund2012]. These junctions support non-chiral Majorana modes [Fu2008, Hasan2010, Qi2011], leading to $4\pi$-periodic Josephson effects [Fu2008, Fu2009, Olund2012, wiedenmann2016].
Dissipation is an intrinsic aspect of real-world Josephson junctions, where non-Hermitian Hamiltonians capture the effects of environmental coupling [cayao2023non, Kornich2023, Li2024, Shen2024, cayao2024non, Beenakker2024]. In non-Hermitian Josephson junctions (NH-JJs), dissipation generates complex spectra and exceptional points (EPs), producing behavior absent in Hermitian systems [Li2024, cayao2023non, Shen2024, Pino2025]. Although EPs in the ABS spectrum may leave no clear imprint on static quantities such as the DC Josephson current [Shen2024], they do manifest in dynamical observables like the current susceptibility [Shen2024] and in thermodynamic measures such as entropy [Pino2025]. However, the impact of non-Hermiticity on topological Josephson junctions remains largely unexplored, particularly in systems where dissipation competes with topological protection.
@@
In this work, we address the central question: how does dissipation caused by coupling a ‘lossy’ metallic lead [cayao2023non, Ohnmacht2024] at the junction, affect the quantum transport properties of TIJJs? To explore this, we model the dissipative effects of the lead by an effective non-Hermitian Hamiltonian [Harrington2022, Li2024] and investigate transport in a non-Hermitian lateral Josephson junction on the surface of a 3DTI (NH-TIJJs), as illustrated schematically in Fig. 1. Here, proximity-induced superconductivity intertwines with non-Hermitian effects, offering a unique platform to study the interplay between dissipation and topological superconductivity.
Effective model of a NH-TIJJ: The NH-TIJJ in Fig. 1 consists of two $s$-wave singlet superconductors with identical gap $\Delta$ and phases $\phi_{L}$ and $\phi_{R}$ on the left and right, respectively. Dissipation due to the ‘lossy’ lead [cayao2023non, Ohnmacht2024] connected to an electron reservoir is modeled in a minimal way, within the Lindblad formalism neglecting quantum jumps [breuer2002, daley2014] (see Appendix A for details), by an imaginary potential barrier at the junction [Li2024]. The combined effect of the top gate and the lead is therefore captured by a complex contact junction potential barrier: $U(x)=(V_{1}-iV_{2})\delta(x)$, where $V_{1}>0$ represents the real barrier strength tunable by the applied gate voltage $V_{g}$ and $V_{2}>0$ accounts for the imaginary barrier strength induced by the lead.
Figure 2: Characteristics of the NH-TIJJ for a complex barrier ($Z_{1}\neq 0$, $Z_{2}\neq 0$) in the tunneling limit ($\Gamma_{0}\ll\Delta$): (a) The real (solid lines) and imaginary (dashed lines) parts of the Andreev spectrum of the NH-TIJJ as a function of the phase difference $\phi$ plotted for different angles of incidence $\theta$ taking the imaginary barrier parameter $Z_{2}=0.2$. (b) Variation of the imaginary part of the spectrum as a function of $\phi$ and $Z_{2}$ for the incident angle $\theta=\pi/3$, the boundary of the dark purple region demonstrates the asymmetric gap-exit points in the spectrum. (c) The zero temperature angle-resolved current $I_{\theta}(\phi)$ plotted as a function of $\phi$ for different angles of incidence. (d) The total supercurrent $I_{\mathrm{total}}(\phi)$ as a function of $\phi$ for different barrier parameters. The current is normalized by $I_{0}=eN\Gamma_{0}/2\hbar$ and we have chosen the real barrier parameter $Z_{1}=0.4$ and tunneling parameter $\gamma=0.3$.
Assuming tunnel coupling between the superconductors and the 3DTI surface state [Tkachov2013, Fu2008, Potter2011, Lababidi2011], the physics of the proximity-induced superconductivity on the surface state can be described by the effective Bogoliubov-de Gennes (BdG) Hamiltonian [Tkachov2013, Fu2008]:
@@
$H_{\rm BdG}$ $=$ $\begin{bmatrix}\frac{\hbar}{i}v_{F}\bm{\sigma}\cdot\nabla-\mu-\Gamma(\epsilon)&\tilde{\Delta}(\epsilon)i\sigma_{y}e^{i\phi_{L,R}}\vskip 2.84544pt\\ -\tilde{\Delta}(\epsilon)i\sigma_{y}e^{-i\phi_{L,R}}&\frac{\hbar}{i}v_{F}\bm{\sigma}^{*}\cdot\nabla+\mu-\Gamma(\epsilon)\end{bmatrix}$,
$\tilde{\Delta}(\epsilon)$ $=$ $\frac{i\Gamma_{0}\Delta}{\sqrt{\epsilon^{2}-\Delta^{2}}}\;,\Gamma(\epsilon)=\frac{i\Gamma_{0}\epsilon}{\sqrt{\epsilon^{2}-\Delta^{2}}}\;,\Gamma_{0}=\pi t^{2}\mathcal{D}_{0};$ (1)
@@
written in the Nambu basis $\Phi^{\dagger}=(\varphi^{\dagger}_{\uparrow},\varphi^{\dagger}_{\downarrow},\varphi_{\uparrow},\varphi_{\downarrow})$ with $\varphi^{\dagger}_{s}$ being the electron creation operator with spin $s=\uparrow,\downarrow$. $\mu$ is the chemical potential of the superconductor, $v_{F}$ is the Fermi velocity of the surface states of the 3DTI and $\sigma_{i}$ ($i=1,2$) are the Pauli matrices in spin space. $\tilde{\Delta}(\epsilon)$ is the proximity induced superconducting pairing potential which is a function of the Bogoliubov quasi-particle energy $\epsilon$, $\Gamma(\epsilon)$ is the shift in energy of the quasi-particles due to tunneling with amplitude $t$, and $\Gamma_{0}$ is the tunneling energy scale that determines the tunneling rate. $\mathcal{D}_{0}$ is the density of states of the superconductors in the normal state. The proximitized superconductivity on the 3DTI surface is determined by the tunneling energy $\Gamma_{0}$, in the tunneling regime ($\Gamma_{0}\ll\Delta$) the induced gap is much smaller than the parent s-wave gap and in the strong proximity limit ($\Gamma_{0}=\Delta$) the pairing potential is equal to that of the parent superconductor. In this work, we investigate the transport properties in both of these regimes to see whether any new feature is observed in the tunneling regime.
@@
To investigate transport in the “short” NH-TIJJ (Fig. 1), assuming the width of the junction much smaller than the coherence length of the superconductors, we consider $\phi_{L}=0$ and $\phi_{R}=\phi$ without loss of generality. In this limit, the physics of the system can be described by [Tkachov2013]:
$\left[H_{\text{BdG}}(\bm{r})+(V_{1}\tau_{z}-iV_{2}\tau_{0})\sigma_{0}\right]\Psi(\bm{r})=\epsilon\Psi(\bm{r})\;,$ (2)
@@
where, $\tau_{z}$ is the third Pauli matrix in the Nambu space, $\sigma_{0}$ is the $2\times 2$ identity matrix and $\Psi(\bm{r})=\left[\psi_{\uparrow,\epsilon}(\bm{r}),\psi_{\downarrow,\epsilon}(\bm{r}),\psi^{*}_{\uparrow,-\epsilon}(\bm{r}),\psi^{*}_{\downarrow,-\epsilon}(\bm{r})\right]^{T}$ is the Nambu spinor wavefunction. Quasiparticles enter the barrier from the left at an angle of incidence $\theta\in[-\pi/2,\pi/2]$ and we assume translational invariance along the $y$-direction. Klein tunneling is insensitive to the precise form of the barrier [Tkachov2013, Tkachov2013a], so the specific shape of the potential does not matter for our analysis; only the influence of non-Hermiticity is relevant.
@@
$\psi_{k_{y}}(0_{-})=[1+iZ_{1}\tau_{z}\sigma_{x}+Z_{2}\tau_{0}\sigma_{x}]\psi_{k_{y}}(0_{+})\;,$ (3)
which leads to an equation in terms of $\mathcal{E}/\tilde{\Delta}$ (see Appendix B):
@@
$4iZ_{2}\cos{(\theta)}\frac{\mathcal{E}}{\tilde{\Delta}}\sqrt{1-\left(\frac{\mathcal{E}}{\tilde{\Delta}}\right)^{2}}=\mathcal{F(\theta,\phi)}-2T_{2}\left(\frac{\mathcal{E}}{\tilde{\Delta}}\right)^{2}\,.$ (4)
Here, $\mathcal{F}(\theta,\phi)=(1+Z^{2}_{1}-Z^{2}_{2})\cos^{2}{(\theta)}\cos{(\phi)}+(Z^{2}_{1}+Z^{2}_{2})(1+\sin^{2}{(\theta)})+\cos^{2}{(\theta)}+2Z_{1}Z_{2}\cos^{2}{(\theta)}\sin{(\phi)}$, $T_{2}=Z_{1}^{2}+Z_{2}^{2}+\cos^{2}{(\theta)}$, $\mathcal{E}(\epsilon)=\epsilon+i\Gamma(\epsilon)$ and $\tilde{\Delta}=\tilde{\Delta}(\epsilon)$. The above equation can be rearranged into a quadratic equation in the variable $y=\left|\frac{\mathcal{E}}{\tilde{\Delta}}\right|^{2}$ as $Ay^{2}+By+C=0$ giving the solutions
$\left|\frac{\mathcal{E}}{\tilde{\Delta}}\right|^{2}=\frac{-B\pm\sqrt{B^{2}-4AC}}{2A}\;,$ (5)
where $A=4Z^{2}_{2}\cos^{2}{(\theta)}-T^{2}_{2}$, $B=\mathcal{F}(\theta,\phi)T_{2}-4Z^{2}_{2}\cos^{2}{(\theta)}$, $C=-\mathcal{F}^{2}(\theta,\phi)/4$. In the case of strong proximity effect, $\mathcal{E}$ and $\tilde{\Delta}$} are independent of the quasiparticles energy $\epsilon$ and the induced gap can be treated as constant. In this limit, Eq. (5) can be interpreted as the bound state solutions for the TIJJ whose induced gap is equal to that of the parent superconductor (see Appendix B for more discussion). Denoting the R.H.S of Eq. (5) as $\mathcal{S}(\theta,\phi)$, we substitute the expressions of $\mathcal{E}$ and $\tilde{\Delta}$ and arrive at the following equation for the bound states in terms of $\alpha=|\frac{\epsilon}{\Delta}|$ in the tunneling regime
$\alpha^{2}(1-\alpha^{2})=\gamma^{2}\left(\sqrt{\mathcal{S}(\theta,\phi)}-\alpha\right)^{2}\;,$ (6)